OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. I WILL GIVE MEDALS

OpenStudy (anonymous):

3(x+1) = -2(x-1) + 6 Distribute the 3 and the -2 3x+3=-2x+2+6 add 2x to both sides. subtract 3 from both sides. 5x= 5 divide both sides by 5 x = 1 that help?